A potter's wheel moves uniformly from rest to an angular speed of 1.00 rev/s in 30.0 s. (a) Find its average angular acceleration in radians per second per second. (b) Would doubling the angular acceleration during the given period have doubled the final angular speed?
1
Expert's answer
2020-10-27T11:15:44-0400

a)


"\alpha=\frac{1(2\pi)}{30}=0.209\frac{rad}{s^2}"

b) Yes.